# Quadratic Gaussian Splatting for Efficient and Detailed Surface Reconstruction

### v1

@@ 3.2.2 Splatting
Ray-splat Intersection. Let the camera center in the Gaussian local space be denoted as $\hat{\mathbf{o}}\in\mathbb{R}^{3\times 1}$ and the ray direction as $\hat{\mathbf{d}}\in\mathbb{R}^{3\times 1}$. A point on the ray can be defined as $\hat{\mathbf{p}}=\hat{\mathbf{o}}+t\hat{\mathbf{d}}$. By substituting $\hat{\mathbf{p}}$ into the Equation 8, we solve a quadratic equation to obtain two intersection points: the nearer point $\hat{\mathbf{p}}_{n}=(\hat{x}_{n},\hat{y}_{n},\hat{z}_{n})$ and the farther point $\hat{\mathbf{p}}_{f}=(\hat{x}_{f},\hat{y}_{f},\hat{z}_{f})$, with $t_{n}\leq t_{f}$. To ensure multi-view consistency in QGS, we select one of these two points: First, if the geodesic distance of $\hat{\mathbf{p}}_{n}$ is within $3\sigma(\theta_{n})$, we choose $\hat{\mathbf{p}}_{n}$. If not, we check if $\hat{\mathbf{p}}_{f}$ is within $3\sigma(\theta_{f})$. If so, we select $\hat{\mathbf{p}}_{f}$. If neither condition is met, we assume no intersection between the ray $\hat{\mathbf{p}}(t)$ and the primitive. The derivation is provided in the supplementary material.
@@
We render the normal map $N(\mathbf{p})$ and curvature map $K(\mathbf{p})$ by Equation 16 for a given viewpoint using alpha-blending, as shown in the Fig 4.
$N(u,v)$ $=\sum_{i=0}^{N-1}G_{i}\alpha_{i}\textbf{n}_{i}\prod_{j=0}^{i-1}(1-G_{j}\alpha_{j})$ (16)
$K(u,v)$ $=\sum_{i=0}^{N-1}G_{i}\alpha_{i}K_{i}\prod_{j=0}^{i-1}(1-G_{j}\alpha_{j})$
normal
curvature
Figure 4: The normal map (left) and curvature map (right) of a scene. In the curvature map, blue indicates higher curvature and curved areas, while red indicates lower curvature and flatter areas.
Per-tile Sorting and Per-pixel Resorting. In Equation 16, the index $i$ is determined by sorting primitives from near to far. Most GS-like methods sort by splat centroid depth rather than ray-splat intersection depth, which, as noted by StopThePop [25], causes popping artifacts. Additionally, we observed that centroid sorting introduces streak artifacts in geometry reconstruction, as shown in Fig 8. To address this, we apply StopThePop’s Per-tile Sorting and Per-pixel Resorting [25] to our QGS. In short, we first compute the ray closest to the quadric center in each tile to estimate intersection depth for rough sorting. During volume rendering, we use a buffer array to locally resort quadrics based on the intersection depth of each pixel. Further details are provided in the supplementary material.
@@ 3.3 Optimization
For QGS, which has stronger geometric fitting capability, relying solely on photometric consistency error can introduce geometric noise, such as unsmooth surfaces. Therefore, we adopt the depth distortion loss and normal consistency loss proposed by 2DGS [12] to mitigate this.
Depth Distortion. We employ the depth distortion loss to promote tighter alignment of the quadric surfaces.
@@
Here, $\omega_{i}=\bar{\alpha}_{i}T_{i}$ denotes the alpha-blending weight of the i-th Gaussian, and $t_{i}$ represents the depth at the ray-splat intersection. $i,j$ index the Gaussians contributing to the ray. Additionally, following GOF’s recommendation, we freeze the gradient of the depth distortion loss through $\omega_{i}$, optimizing only the depth $t_{i}$.
Normal Consistency. 2DGS [12] introduces the normal consistency loss to ensure that all primitives on a ray are locally aligned with the actual surface.
@@
Here, $\mathbf{n}_{i}$ represents the splat normal facing the camera, while $\mathbf{N}$ is computed by differentiating the depth point $\mathbf{p}$ from neighboring pixels.
$\mathbf{N}(u,v)=\frac{\nabla_{u}\mathbf{p}\times\nabla_{v}\mathbf{p}}{|\nabla_{u}\mathbf{p}\times\nabla_{v}\mathbf{p}|}$ (19)
However, neighboring pixels may not satisfy the local planar assumption, especially in regions with significant depth variation. Using differential normal consistency loss in edge areas can introduce errors, contributing to the over-smoothing observed in 2DGS. PGSR [4] and MVG-splatting [19] have both noted this issue and use image edges to approximate geometric edges, applying additional processing at these edges. We have found that, in most scenes, image edges do not fully correspond to geometric edges, especially in uniformly lit areas. Thus, we utilize the curvature map, which more accurately corresponds to geometric edges and is both efficiently and uniquely generated by QGS, to guide normal supervision.
@@
$\mathcal{L}=\mathcal{L}_{c}+\lambda_{d}\mathcal{L}_{d}+\lambda_{n}\mathcal{L}_{Kn}$ (21)
@@ 4 Experiment
Figure 5: Qualitative geometric reconstruction comparisons on three datasets (DTU, Tanks and Temples, and Mip-NeRF 360) against 2DGS and GOF, covering indoor, outdoor, object-centric, and forward-facing scenes. QGS shows obvious improvements in both detail and smoothness compared to 2DGS and GOF.
CD (mm)$\downarrow$ 24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean Time
NeuS [30] 1.00 1.37 0.93 0.43 1.10 0.65 0.57 1.48 1.09 0.83 0.52 1.20 0.35 0.49 0.54 0.84 $>$12h
VolSDF [32] 1.14 1.26 0.81 0.49 1.25 0.70 0.72 1.29 1.18 0.70 0.66 1.08 0.42 0.61 0.55 0.86 $>$12h
Neuralangelo [18] 0.37 0.72 0.35 0.35 0.87 0.54 0.53 1.29 0.97 0.73 0.47 0.74 0.32 0.41 0.43 0.61 $>$128h
@@
Ours 0.46 0.81 0.40 0.42 1.07 0.80 0.79 1.25 0.95 0.67 0.62 1.27 0.40 0.60 0.49 0.73 0.83h
Table 1: Quantitative Chamfer Distance comparison of QGS with GS-like and NeRF-like methods on the DTU dataset. Compared to the planar representation in 2DGS, the surface-based QGS shows a clear improvement in geometric reconstruction quality. Additionally, QGS is an order of magnitude faster than NeRF-like methods.
F1-Score $\uparrow$ NeuS Geo-NeuS Neuralangelo 2DGS GOF Ours
Barn 0.29 0.33 0.70 0.41 0.51 0.43
Caterpillar 0.29 0.26 0.36 0.24 0.41 0.31
Courthouse 0.17 0.12 0.28 0.16 0.28 0.26
Ignatius 0.83 0.72 0.89 0.52 0.68 0.79
Meetingroom 0.24 0.20 0.32 0.17 0.28 0.25
Truck 0.45 0.45 0.48 0.45 0.58 0.60
Mean 0.38 0.35 0.50 0.33 0.46 0.44
Time $>$24h $>$24h $>$127h 0.57h 2h 2h
Table 2: Quantitative F1-Score comparison of QGS with GS-like and NeRF-like methods on the TNT dataset. QGS shows notable gains over 2DGS and is competitive with prior SOTA methods.
@@
Rasterizer We implemented Quadratic Splatting with custom CUDA kernels on the 3DGS framework, extending the renderer to output depth distortion maps, depth maps, normal maps, and curvature maps. Given the non-convex nature of paraboloids and the lack of an inverse for the geodesic distance function, we used rectangular truncation and approximations to compute image bounding boxes, as detailed in the supplementary materials. To handle the more complex surface intersections, we incorporate Per-tile Sorting, Per-pixel Local Resorting from StopThePop [25], with further details also in the supplementary materials.
@@
Settings. We adopted the adaptive control strategy from 3DGS [16]. Similar to 2DGS [12], the QGS splatting process approximates by projecting the 3D center gradient onto screen space instead of relying directly on the 2D projected center gradient, using a gradient threshold of 0.3 and a percent dense value of 0.001. To ensure a similar number of Gaussian primitives across methods, we also set the gradient threshold for other methods to 2e-4. All experiments were conducted on a single A6000 GPU.
@@
Mesh Extraction. We used a weighted combination of 0.5 median depth (i.e., $t_{\mathrm{median}}=\max{t_{i}|T_{i}>0.5}$) and 0.5 expected depth (i.e., $t_{\mathrm{expected}}=\sum_{i=0}^{N-1}G_{i}\alpha_{i}T_{i}t_{i}$) as the final output depth. Then we fused the depth maps using Truncated Signed Distance Fusion (TSDF) with Open3D [37]. Following the 2DGS setup, we set the voxel size to 0.004 and truncation threshold to 0.002 for fair comparison.
4.2 Comparison
Geometry Evaluation In all experiments, we performed evaluations at half image resolution. In Table 1, we compared QGS with implicit [30, 32, 18] and explicit [12, 35, 11] SOTA reconstruction methods on the DTU dataset using Chamfer Distance as the benchmark. QGS outperformed all explicit methods, and while a gap remains with Neuralangelo [18], our reconstruction time is two orders of magnitude faster. In certain scenes, such as Scan 97 and Scan 105, QGS even outperformed Neuralangelo. In Table 2, we evaluated QGS and other SOTA methods on the TNT dataset using F1-score. In large-scale scenes, QGS consistently outperformed 2DGS [12]. Notably, in the Courthouse and Meetingroom scenes, QGS captured more details due to its higher geometric flexibility, as shown in the Fig 6.
2DGS
QGS
@@ 4.2 Comparison
Figure 6: Mesh reconstruction comparison between QGS and 2DGS in large-scale scenes. The first row shows the Courthouse scene, with both using median depth, and the second row shows the Meetingroom scene, using a 0.5 weighted combination of median and expected depth.
As shown in Fig 5, our QGS achieves more complete geometry and stronger detail capture due to the greater flexibility of individual primitives. Note that our mesh extraction is limited by the lower resolution of the voxel grid in TSDF fusion, resulting in slightly lower accuracy compared to GOF [35], which uses marching tetrahedra, leading to longer meshing times and higher VRAM consumption.
QGS’s second-order fit greatly improves on 2DGS [12], capturing finer details and providing second-order geometric information. The curvature map in Fig 4 illustrates how QGS better conforms to object edges with higher curvature, while adhering to flatter surfaces in other areas. It is important to note that the advantage of surface-based representations lies in their ability to provide multi-view consistent depth and normals (and curvature in our method). Thus, both theoretically and practically, it is straightforward to transfer multi-view geometric techniques, as done in PGSR [4] and MVG-Splatting [19], to QGS. We leave this exploration for future work to further enhance our approach.
@@
Rendering Evaluation We compared rendering quality on the Mip-NeRF360 dataset with baseline approaches. As a surface-based method, QGS achieves competitive rendering results in indoor scenes, as shown in Table 3, but performs less effectively in outdoor environments. We attribute this to QGS’s higher geometric fitting capability, which may lead to overfitting in regions with sparse views or low texture. Future work could explore additional regularization constraints, particularly on the curvature of quadric surfaces, to address this issue.
@@ 4.3 Ablation
Curvature-Guided Normal Consistency Loss. As mentioned in Section 3.2, we use curvature to guide normal consistency supervision. High-curvature regions are treated as high-frequency areas that violate the local planar assumption, so we reduce supervision in these regions. As shown in Table 4 and Fig 7, curvature-guided normal supervision helps recover more scene details, leading to improved geometric quality.
w/o $\lambda_{K}$
w/ $\lambda_{K}$
Figure 7: Comparison of curvature-guided normal consistency versus without curvature guidance. Curvature-guided supervision effectively reduces over-smoothing and recovers more details.
F1-Score $\uparrow$ Truck Ignatius
@@
w/o $\lambda_{K}$ 0.63 0.51 0.57 0.74 0.75 0.75
w/ $\lambda_{K}$ 0.65 0.55 0.60 0.78 0.80 0.79
Table 4: Quantitative F1-Score comparison on the TNT dataset with and without curvature-guided normal supervision.
Per-tile Sorting. 2DGS suggests the actual surface lies at the median intersection point where opacity reaches 0.5. However, most GS-like methods sort Gaussians by centroid depth, leading to streak-like inconsistencies, as shown in Fig 8. To address this, we implemented per-tile sorting (TS) and per-pixel resorting (PR) based on the methods proposed in StopThePop [25], further refining them to remove inconsistencies and enhance reconstruction quality, as shown in Table 5.
w/o TS and PR
w/ TS
w/ TS and PR
Figure 8: Normal comparison with and without per-tile sorting. Centroid depth sorting causes noticeable streak-like inconsistencies, also seen in 2DGS.
CD (mm)$\downarrow$ 24 37 97 122
w/o TS and PR 0.48 0.85 1.04 0.55
w/ TS 0.46 0.80 0.96 0.52
w/ TS + PR 0.46 0.81 0.95 0.49
Table 5: Quantitative Chamfer Distance comparison on the DTU dataset with and without per-tile sorting.

### v3

@@ 3.2.1 Splatting
Ray-splat Intersection. Let the camera center in the Gaussian local space be denoted as $\hat{\mathbf{o}}\in\mathbb{R}^{3\times 1}$ and the ray direction as $\hat{\mathbf{d}}\in\mathbb{R}^{3\times 1}$. A point on the ray can be defined as $\hat{\mathbf{p}}=\hat{\mathbf{o}}+t\hat{\mathbf{d}}$. By substituting $\hat{\mathbf{p}}$ into the Equation 6, we solve a quadratic equation to obtain two intersection points: the nearer point $\hat{\mathbf{p}}_{n}=(\hat{x}_{n},\hat{y}_{n},\hat{z}_{n})$ and the farther point $\hat{\mathbf{p}}_{f}=(\hat{x}_{f},\hat{y}_{f},\hat{z}_{f})$, with $t_{n}\leq t_{f}$. To ensure multi-view consistency in QGS, only one valid point of the two is selected. First, if the geodesic distance of $\hat{\mathbf{p}}_{n}$ is within $3\sigma(\theta_{n})$, we choose $\hat{\mathbf{p}}_{n}$. If not, we check if $\hat{\mathbf{p}}_{f}$ is within $3\sigma(\theta_{f})$. If so, we select $\hat{\mathbf{p}}_{f}$. If neither condition is met, we assume no intersection between the ray $\hat{\mathbf{p}}(t)$ and the primitive. This assumption is generally valid because the significant weight typically causes the first point to occlude the second point. The derivation is provided in the supplementary material.
@@
We render the normal map $\mathbf{N}$ and curvature map $\mathbf{K}$ by Equation 14 for a given viewpoint using alpha-blending, as shown in the Fig 6.
$\{\mathbf{N},\mathbf{K}\}$ $=\sum_{i=0}^{N-1}G_{i}\alpha_{i}\prod_{j=0}^{i-1}(1-G_{j}\alpha_{j})\{\textbf{n}_{i},K_{i}\}$ (14)
(a) w/o resorting
(b) w/ resorting
Figure 7: Comparison of depth normals with and without resorting. Centroid-only sorting (a) introduces stripe artifacts, while resorting (b) produces smoother and more consistent normals.
Per-pixel Resorting. We adopt the per-pixel resorting method from Stopthepop to address the depth pop-out issue, as illustrated in Fig 7. Implementation and derivation details are provided in the supplementary material.
@@ 3.3 Optimization
We now introduce the our training objectives during optimization.
Depth Distortion. We adopt the depth distortion loss and normal consistency loss proposed by 2DGS [16].
@@
Here, $\omega_{i}=\bar{\alpha}_{i}T_{i}$ denotes the alpha-blending weight of the i-th Gaussian, and $t_{i}$ represents the depth at the ray-splat intersection. $i,j$ index the Gaussians along the ray.
Curvature Guided Normal Consistency. 2DGS [16] introduces the normal consistency loss to ensure that all primitives on a ray are locally aligned with the actual surface.
@@
Here, $\mathbf{n}_{i}$ represents the splat normal facing the camera, while $\mathbf{N}$ is computed by differentiating the depth point $\mathbf{p}$ from neighboring pixels. However, neighboring pixels may violate the local planar assumption, particularly in edge regions. Using normal consistency loss in edge areas can introduce errors, contributing to the over-smoothing observed in 2DGS. A straightforward approach is to approximate geometric edges using image edges, which can then guide normal consistency supervision [5, 25]. However, we found that, in most scenes, image edges do not fully correspond to geometric edges, especially in uniformly lit areas. Thus, we use the curvature map, which more accurately corresponds to geometric edges and is efficiently and uniquely generated by QGS, to guide normal supervision.
@@
Multi-view regularization Loss. To ensure a fairer comparison with plane-based PGSR [5], we incorporate PGSR’s multi-view regularization, computing warp and photometric consistency losses via homography transformations.
$\mathbf{H}_{rn}=$ $\mathbf{K}_{n}(\mathbf{R}_{rn}+\frac{\mathbf{T}_{rn}\mathbf{n}_{r}^{T}}{\mathbf{n}_{r}^{T}\mathbf{X}_{r}})\mathbf{K}_{r}^{-1}$ (18)
[…]
$(1-NCC(\mathbf{I}_{r}(\mathbf{p}_{r}),\mathbf{I}_{n}(\mathbf{H}_{rn}\mathbf{p}_{r})))$
Here, $\mathbf{n}_{r}$ represents the normal map, and $\mathbf{X}_{r}$ denotes the 3D points projected from the depth map. $\mathbf{R}_{rn}$ represents the relative rotation, $\mathbf{K}$ is the camera intrinsic, $\mathbf{p}_{r}$ denotes the pixel coordinates, and $\mathbf{I}$ refers to the image intensity. This loss enforces multi-view constraints by computing the reprojection error from the depth and normal maps, along with normalized cross-correlation (NCC) [47].
@@
$\mathcal{L}=\mathcal{L}_{c}+\lambda_{d}\mathcal{L}_{d}+\lambda_{n}\mathcal{L}_{Kn}+\lambda_{Mv}\mathcal{L}_{Mv}$ (19)
@@ 4 Experiment
scan24
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
scan65
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
scan118
Reference Image
Refer to captionRefer to caption
2DGS
Refer to captionRefer to caption
GOF
Refer to captionRefer to caption
PGSR
Refer to captionRefer to caption
Ours
Figure 8: Qualitative geometric reconstruction comparisons on the DTU dataset. Our method achieves reconstructions of higher quality and greater detail.
[…]
NeuS [42] 1.00 1.37 0.93 0.43 1.10 0.65 0.57 1.48 1.09 0.83 0.52 1.20 0.35 0.49 0.54 0.84 >>12h
VolSDF [45] 1.14 1.26 0.81 0.49 1.25 0.70 0.72 1.29 1.18 0.70 0.66 1.08 0.42 0.61 0.55 0.86 >>12h
Neuralangelo [24] 0.37 0.72 0.35 0.35 0.87 0.54 0.53 1.29 0.97 0.73 0.47 0.74 0.32 0.41 0.43 0.61 >>128h
3DGS [21] 2.14 1.53 2.08 1.68 3.49 2.21 1.43 2.07 2.22 1.75 1.79 2.55 1.53 1.52 1.50 1.96 11.2min
Gaussian surfels [8] 0.66 0.93 0.54 0.41 1.06 1.14 0.85 1.29 1.53 0.79 0.82 1.58 0.45 0.66 0.53 0.88 6.7min
@@
GSDF [48] 0.59 0.94 0.46 0.38 1.30 0.77 0.73 1.59 1.29 0.76 0.59 1.22 0.38 0.52 0.51 0.80 32min
GS-pull [52] 0.51 0.56 0.46 0.39 0.82 0.67 0.85 1.37 1.25 0.73 0.54 1.39 0.35 0.88 0.42 0.75 22min
Ours w/o MV 0.46 0.76 0.40 0.38 0.92 0.80 0.76 1.25 0.95 0.67 0.62 1.20 0.38 0.60 0.47 0.71 25min
PGSR [5] 0.40 0.60 0.39 0.37 0.78 0.59 0.53 1.18 0.67 0.63 0.48 0.62 0.34 0.42 0.39 0.56 40min
Ours 0.38 0.62 0.37 0.38 0.75 0.55 0.51 1.12 0.68 0.61 0.46 0.58 0.35 0.41 0.40 0.54 48min
Table 1: Quantitative Chamfer Distance comparison on the DTU dataset. Our method achieves the best performance both with and without multi-view regularization.  ,  ,   denote the best, second best, and third best results, respectively.
F1-Score $\uparrow$ Geo-NeuS N-angelo 2DGS GOF Ours w/o MV PGSR Ours
Barn 0.33 0.70 0.41 0.51 0.46 0.52 0.55
Caterpillar 0.26 0.36 0.24 0.41 0.32 0.38 0.40
Courthouse 0.12 0.28 0.16 0.28 0.26 0.26 0.28
Ignatius 0.72 0.89 0.52 0.68 0.79 0.77 0.81
Meetingroom 0.20 0.32 0.17 0.28 0.25 0.29 0.31
Truck 0.45 0.48 0.45 0.58 0.60 0.62 0.64
Mean 0.35 0.50 0.33 0.46 0.45 0.47 0.50
Time >>24h >>127h 34min 114min 43min 66min 75min
Table 2: Quantitative F1-Score comparison of QGS with GS-like and NeRF-like methods on the TNT dataset. QGS outperforms all methods, achieving state-of-the-art reconstruction results.
@@
Rasterizer. We implemented Quadratic Splatting with custom CUDA kernels on the 3DGS framework [21], extending the renderer to output depth distortion maps, depth maps, normal maps, and curvature maps. Since quadrics can be non-convex, we used rectangular truncation and approximations to compute image bounding boxes, as detailed in the supplementary materials.
@@
Settings. We adopted the adaptive control strategy from 3DGS [21]. Similar to 2DGS [16], QGS projects the 3D center gradient onto screen space instead of using the 2D projected gradient. A gradient threshold of 0.3 and a percent dense value of 0.001 ensure consistent point cloud number with other methods. All experiments were conducted on a single A6000 GPU.
@@
Mesh Extraction. We used median depth (i.e., $t_{\mathrm{median}}=\max{t_{i}|T_{i}>0.5}$) as the final output depth. Then we fused the depth maps using Truncated Signed Distance Fusion (TSDF) with Open3D [53]. Following the 2DGS setup, we set the voxel size to 0.004 and truncation threshold to 0.02.
4.2 Comparison.
Geometry Evaluation. In all experiments, evaluations were performed at half the image resolution. To clearly demonstrate the improvement introduced by our quadric surfels and to maintain consistency with other methods, we denote the variant of QGS without multi-view regularization loss as QGS w/o MV. In Table 1, we compared QGS with implicit [42, 45, 24] and explicit [16, 50, 13, 48, 5, 52] SOTA reconstruction methods on the DTU dataset using Chamfer Distance as the metric. As shown in Table 1, our method outperforms both implicit and explicit approaches on the DTU dataset. Furthermore, even without multi-view regularization, our method surpasses others in accuracy while maintaining competitive speed. This improvement stems from the superior geometric fitting capability of quadrics, which enables finer detail preservation, as illustrated in Fig 8.
In Table 2, we also evaluated QGS and other SOTA methods [11, 24, 16, 50, 5] on the TNT dataset using F1-score. Without multi-view regularization, our method significantly outperforms the disk-based 2DGS and achieves comparable results to GOF while requiring only half the computation time. With multi-view regularization, it surpasses all methods in reconstruction accuracy. Additionally, we provide a qualitative comparison between QGS without multi-view regularization and 2DGS to further demonstrate the enhanced geometric fitting capability of quadric surfels, as shown in Fig 9.
(a) 2DGS
(b) QGS w/o MV
@@
Figure 9: Mesh reconstruction comparison between 2DGS and QGS w/o MV. The first row shows the Meetingroom in TNT dataset, and the second row shows the kitchen scene in Mip-NeRF 360 dataset.
@@
Rendering Evaluation. We compared rendering quality on the Mip-NeRF360 dataset with baseline approaches. As a surface-based method, QGS achieves competitive rendering results in indoor scenes, as shown in Table 3, but performs less effectively in outdoor environments. We attribute this to QGS’s higher geometric fitting capability, which may lead to overfitting in regions with sparse views or low texture. Future work could explore additional regularization constraints, particularly on the curvature of quadric surfaces, to address this issue.
@@ 4.3 Ablation
In this section, we assess the impact of individual QGS components on reconstruction quality, including per-pixel resorting, curvature-guided normal consistency, and multi-view regularization, as presented in Table 4 and Fig 10. We denote w/o sort as the absence of per-pixel resorting, w/o $\lambda_{K}$ as the absence of curvature-guided normal consistency, and w/o MV as the absence of multi-view regularization. We observe that: (a) Disabling curvature guidance fills fine gaps in the normal map, reducing reconstruction quality. (b) The absence of per-pixel sorting similarly fills gaps and leads to uneven surfaces. (c) Disabling multi-view regularization causes overfitting in low-texture regions, resulting in surface artifacts, such as dents on the front of the truck.
Additionally, we conducted ablation studies without multi-view regularization to isolate and evaluate the effectiveness of quadrics. Specifically, we fixed the quadric’s z-axis scale to 0.001, approximating it as a disk (denoted as $s_{3}$ fixed). We also evaluated optimization using Euclidean distance (denoted as w/ Euclid). Both modifications significantly reduced reconstruction quality, as shown in Table 5.
F1-Score $\uparrow$ B C CH I M T mean
Full model 0.55 0.40 0.28 0.81 0.31 0.64 0.50
w/o Sort 0.53 0.39 0.26 0.77 0.29 0.62 0.48
w/o $\lambda_{K}$ 0.56 0.37 0.27 0.80 0.29 0.62 0.49
w/o MV 0.46 0.32 0.26 0.79 0.25 0.60 0.45
Table 4: Quantitative ablation study on the TNT dataset. We present ablation results for Barn (B), Caterpillar (C), Courthouse (CH), Ignatius (I), Meeting Room (M), and Truck (T). Our full setting achieves the best performance.
(a) w/o $\lambda_{K}$
(b) w/o Sort
(c) w/o MV
(d) Full
Figure 10: Ablation study. The full setting achieves the highest reconstruction quality.
Setting w/o MV $s_{3}$ fixed w/ Euclid
@@
Table 5: Ablation study without multi-view regularization on the TNT dataset. Fixing the quadric as a disk or using Euclidean distance optimization both result in reduced reconstruction quality.